OpenStudy (anonymous):

vertical and horizontal asymptotes of 1/(X-2)+10 How do you tell which the asymptotes apart, I know the answers are 2 and 10 but which one is the V.A and the H.A?

OpenStudy (anonymous):

x in bottom is a clue of vertical asymptote

OpenStudy (anonymous):

so the V.A: is 2 then?

OpenStudy (anonymous):

Well, how did you find 2? You cheated on the calculator. :) Here is an easy method. The minute you see x in the bottom your alarm should go off. x can not be 0 in the denominator so set what ever in bottom to 0; x-2=0, x=2 x cannot be 2 vertical asymptote. Square root should also ring the alarm.

OpenStudy (anonymous):

i actually didn't use a calculator....ummm I know it is 2 because if 1/(2-2) = 1/0 which is undefined an then I know it is 10 because it is a number by itself and I know the 10 will affect the graph by moving the graph up 10 spaces I just don't know which one is which

OpenStudy (anonymous):

Zero in the denominator vertical asymptote. It is going to be second nature. From this point on for the rest of your life, you will remember it.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Here is a term, you can show off mathematical terminology. They like to say the graph blows up.